# Measure Twice, Match Once: Flexible Graph Similarity Computation With A Proactive Optimization Strategy

### v1

@@ 4.2. Node Matching Cost Estimation
(7) $\begin{gathered}\text{dist}(s,t)=\mathbf{h}_{s}-\mathbf{h}_{t}\\ c(s,t)=\text{Estimator}\bigg{(}\text{Concat}(cost_{(s,t)},\text{dist}(s,t))\bigg{)}\end{gathered}$
The input cost $cost_{(s,t)}\in\mathbb{R}^{|op|}$ represents predefined costs per operation type between $s$ and $t$. The concatenation of $cost_{(\cdot,\cdot)}$ and $\text{dist}(\cdot,\cdot)$ is fed into the $\text{Estimator}(\cdot)$, an MLP, which consists of fully connected layers, meaning all input entries are connected to each output. This allows the operation costs and the corrections made by the corresponding operations to be combined in the outputs through learnable weights and biases. Additionally, this $\text{Estimator}(\cdot)$ also serves as a compensation function $g(\cdot)$ to adjust the node matching costs for duplicated edge costs by decoding the neighborhood interactions captured within the node embeddings. While more sophisticated compensation methods could potentially enhance accuracy, this has neither been discussed nor explored in the learning-based GED computation literature. In this work, we prioritize efficiency and consider it an open area for future research. Gen is the first work to consider operation costs before the matching process, correcting the misinterpretation of the role of operation costs in previous methods. This simple solution ensures that, while inherent differences between compared elements remain consistent, they can be dynamically adjusted to accommodate varying cost settings. It is theoretically in line with the principle of GED computation (Eq. 1), enhancing the model’s understanding of GED computation. Experiment 5.2.3 further highlights the limitations of existing methods on this issue and validates the effectiveness of our solution, enabling flexible graph similarity computation.
4.3. Proactive Optimization Strategy
Unlike previous work that relies on the insufficient node-level guidance of matching costs to reactively refine $\mathbf{1}_{s,t}$ in Eq. 5. Gen proposes a proactive optimization strategy that offers guidance from a graph perspective to minimize the need for refinements. Gen first computes the global cost ratio for each match, then uses the total cost ratios of each node’s associated matches to compute the node’s alignment difficulty, which serves as initial guidance. Next, Gen incorporates the interdependencies of matches within and between graphs through a difficulty propagation mechanism.
4.3.1. Guidance Initialization
Global Cost Ratio Computation.
@@ 4.3.1. Guidance Initialization
Matching costs can vary significantly across matches. While the absolute cost of each match may be clear, comparing these costs directly can be misleading, as some matches might seem optimal for one node but not contribute optimally to the overall graph alignment, leading to myopic, locally optimal decisions for individual nodes. To address this, Gen proposes first normalizes matching costs across all potential node matches as global cost ratios, enabling global comparability:
(8) $\gamma(s,t)=\frac{exp(c(s,t))}{\sum_{s\in\mathcal{V}_{S},t\in\mathcal{V}_{T}}exp(c(s,t))}$
@@
Where $\gamma(s,t)\in\mathbb{R}^{d_{out}}$ represents the cost ratio of a match compared with the overall cost of all candidate matches. Intuitively, a match with a higher matching cost can have a larger ratio, reflecting its global significance within the graph-level alignment.
Node Alignment Difficulty Computation.
Each node is involved in different candidate matches, forming a spectrum of differences between the node and the other graph. Gen gathers the cost ratios for each involved source and target node to compute the node alignment difficulty in the global context as follows.
(9) $\tilde{\gamma}(s)=\sum_{i\in\mathcal{V}_{T}}\gamma(s,i)\quad\text{and}\quad\tilde{\gamma}(t)=\sum_{i\in\mathcal{V}_{S}}\gamma(i,t)$
@@
Intuitively, the higher $\tilde{\gamma}(\cdot)$ means the node is involved in high-cost matches, suggesting that accurate matching of this node is crucial, or it may contribute more to the overall matching cost. This guidance provides a comprehensive assessment of each node’s relative importance across the graphs, helping prioritize nodes that are more critical for global alignment.
4.3.2. Difficulty Propagation
Intra-graph Consideration
Matches within a graph are interdependent due to node connections. Matching one node can constrain the possible matches for other nodes, as the neighborhood structure must be preserved to minimize differences. Similarly, the alignment difficulties of neighboring nodes are important. A node with high alignment difficulty can impose constraints on its neighbors, necessitating careful handling of their alignment. To capture such intra-graph relationships, Gen respects the node connections within the graphs and considers the alignment difficulty of each node’s neighborhood. This broadens the decision-making scope from nodes to neighborhoods. Specifically, Gen propagates the alignment difficulty of each node within its own graph as follows.
(10) $\hat{\gamma}(\mathcal{N}(v))=\text{MLP}\bigg{(}\text{Sum}(\tilde{\gamma}(v),\text{Sum}(\tilde{\gamma}(u):\mathcal{A}(v,u)=1))\bigg{)}$
Here, $\hat{\gamma}(\mathcal{N}(v))\in\mathbb{R}^{d_{out}}$ is the raw neighborhood alignment difficulty of each node, where $v\in(\mathcal{V}_{S}\cup\mathcal{V}_{T})$. $\mathcal{A}$ is the adjacency matrix of their respective graphs. We further combine the aggregated difficulties with representations $\mathbf{h}_{v}$, where $\mathbf{h}_{v}$ serve as residual information to strengthen the connections between difficulties and the corresponding representations, as follows.
(11) $\gamma(\mathcal{N}(v))=\text{MLP}\bigg{(}\text{Concat}(\hat{\gamma}(\mathcal{N}(v)),\mathbf{h}_{v})\bigg{)}$
@@ 4.3.2. Difficulty Propagation
Where $\gamma(\mathcal{N}(v))\in\mathbb{R}^{d_{out}}$ is the neighborhood alignment difficulty for each node $v\in(\mathcal{V}_{S}\cup\mathcal{V}_{T})$.
Inter-graph Consideration and Certainty Computation.
@@
A match connects two nodes within different graphs, each node can have a distinct neighborhood alignment difficulty. In this step, Gen considers such inter-graph relationships to evaluate the certainty for each match, balancing the neighborhood alignment difficulties of the involved nodes. Specifically, it concatenates the neighborhood alignment difficulties and leverages a MLP to compute the matching certainty, approximating $\mathbf{1}_{s,t}$ as follows.
(12) $\mathbf{1}_{s,t}=\text{MLP}\bigg{(}\text{Concat}(\gamma(s),\gamma(t))\bigg{)}$
Where $\mathbf{1}_{s,t}\in\mathbb{R}^{d_{out}}$ represents the certainty of each match, enabling more informed decisions by considering the interdependencies between matches.
4.4. Prediction and Training Objective
Predictions. Based on the certainty of each match, Gen softly selects optimal matches in a single step for each node, and the matching cost for each node is calculated as follows.
(13) $c(s)=\sum_{i\in\mathcal{V}_{T}}\mathbf{1}_{s,i}\cdot c(s,i)\quad\text{and}\quad c(t)=\sum_{i\in\mathcal{V}_{s}}\mathbf{1}_{i,t}\cdot c(i,t)$
Finally, the total matching cost of each graph is calculated as the sum of certainty-weighted node matching costs, according to Eq. 5. Then, the GED is computed based on the average of the overall matching costs of the input graph pair.
(14) $\begin{gathered}c(S,T)_{S}=\sum_{s\in\mathcal{V}_{S}}c(s)\quad\text{and}\quad c(S,T)_{T}=\sum_{t\in\mathcal{V}_{S}}c(t)\\ \text{GED}(S,T)=\text{Linear}\bigg{(}\text{Mean}(c(S),c(T))\bigg{)}\end{gathered}$
@@ 4.4. Prediction and Training Objective
Where $c(S,T)_{S}$ and $c(S,T)_{T}$ are the total costs of mappings established for $S$ and $T$, respectively. $\text{GED}(S,T)\in\mathbb{R}$ is the predicted GED, $\text{Linear}(\cdot)$ reduces the dimensionality from $d$ to $1$. This process indirectly encourages consistent mappings between the $S$ and $T$, or the average matching cost increases accordingly.
Training Objective. Given the GED values as supervision signals, which span from 0 to potentially large numbers, we adopt Huber loss (Huber, 1964) for Gen and baselines to enable stable training. It combines the properties of Mean Squared Error (MSE) and Mean Absolute Error (MAE) losses to handle both small and large errors. Given the predicted GED $\hat{y}$ and the ground-truth GED $y$, the loss is computed as follows.
@@
(15) $\mathcal{L}(\hat{y},y)=\begin{cases}\frac{1}{2}(y-\hat{y})^{2}&\text{for }|y-\hat{y}|\leq\delta,\\ \delta(|y-\hat{y}|-\frac{1}{2}\delta)&\text{otherwise. }\end{cases}$
Here, $\delta$ is the threshold for the loss function transition. For errors smaller than $\delta$, it behaves like MSE loss, offering a smooth gradient. For larger errors, it switches to MAE loss, preventing large gradients and improving training stability.
4.5. Complexity Analysis
The complexity of GNN Encoder is $\mathcal{O}(k\cdot|\mathcal{E}|\cdot d_{hidden})$, where $k$ is the number of layers, $\mathcal{E}$ is the number of edges in the input graph, $d_{hidden}$ is the hidden dimensionality during message-passing, we further reduce it to $d_{out}$, and the generated node representations can be reused for each graph as they are pair-independent. The complexity of the Node Matching Cost Estimation is $\mathcal{O}(\text{Max}(|\mathcal{V}_{S}|,|\mathcal{V}_{T}|)^{2}\cdot d_{out})$, where $|\mathcal{V}_{S}|$ and $|\mathcal{V}_{T}|$ are the numbers of nodes in the source and target graphs, respectively. The complexities of Proactive Optimization Strategy is $\mathcal{O}(2\cdot(\text{Max}(|\mathcal{V}_{S}|,|\mathcal{V}_{T}|)^{2}\cdot d_{out}+|\mathcal{E}|\cdot d_{out}))$.
5. Evaluations
In this section, we evaluate Gen against twelve baselines, including two conventional inexact algorithms, on three commonly used GED datasets and synthetic graphs, focusing on: (1) Effectiveness: How accurate is Gen compared with baselines in the GED computation and graph retrieval under a uniform cost setting; (2) Time Efficiency: How efficient is Gen during training and inference; (3) Adaptation to Varying Cost Settings: Can Gen and baselines accommodate non-uniform and varying cost settings; (4) Ablation Study: How effective are our proposed strategy; (5) Hyperparameter Senstitivity: How sensitive is Gen to the number of layers, dimensionality, and GNN backbones; (6) Performance on Larger Synthetic Graphs: How do the models perform on larger synthetic graphs.
5.1. Experimental Setup
5.1.1. Dataset
We conduct experiments on three commonly used GED datasets: AIDS700nef, Linux and IMDBMulti. The graphs within these datasets and the GED values between graph pairs under a uniform cost setting are provided by Bai et al. (2019), resulting in a rich set of graph pairs. The exact GED is computed with A* (Abu-Aisheh et al., 2015) for AIDS700nef and LINUX. Since the exact GED computation is intractable for IMDBMulti, the smallest distances computed by inexact algorithms such as Beam (Neuhaus et al., 2006), VJ (Fankhauser et al., 2011), and Hungarian (Riesen and Bunke, 2009) are adopted as the training targets. For AIDS700nef, we compute GEDs under two non-uniform settings using A* (Abu-Aisheh et al., 2015). Additionally, we generate synthetic graphs, among Erdős-Rényi, power-law, and Watts–Strogatz small-world graphs with a node size ranging from 100 to 400, as extreme data to to assess model performance on larger graphs. Each synthetic graph’s target graph is created by applying $\Delta$ distinct operations, with the number of operations ranging from 10-30% of the edge count. Following prior work (Li et al., 2019; Bai and Zhao, 2021; Piao et al., 2023), $\Delta$ serves as the ground-truth GED for synthetic pairs. We generate 51,200 graphs for training and another 38,400 graphs in three node size groups: $[100,200]$, $[200,300]$, and $[300,400]$, for testing. The statistics of graphs can be found in Table 1.
Table 1. Statistics of datasets.
Dataset #Graphs #Pairs Avg. $|\mathcal{V}|$ Avg. $|\mathcal{E}|$ Max. $|\mathcal{V}|$ Max. $|\mathcal{E}|$
AIDS700nef 700 490K 8.90 8.80 10 14
LINUX 1000 1M 7.58 6.94 10 13
IMDBMulti 1500 2.25M 13.00 65.94 89 1467
Synthetic 51,200 25,6 K 230.24 1789.02 400 8552
5.1.2. Baselines
We includes representative and state-of-the-art GED computation models as baselines: (1) Graph Comparison Methods: GMN-emb (Li et al., 2019), TaGSim (Bai and Zhao, 2021), Greed (Ranjan et al., 2022) and Eric (Zhuo and Tan, 2022); (2) Node Comparison Method: GMN-match (Li et al., 2019), SimGNN (Bai et al., 2019), GraphSim (Bai et al., 2020), GOTSim (Doan et al., 2021), GEDGNN (Piao et al., 2023), GraphEdX (Jain et al., 2024). Additionally, we compare with well-known approximate algorithms, such as the Volgenant–Jonker (VJ) algorithm (Fankhauser et al., 2011) and Beam (Neuhaus et al., 2006). The VJ algorithm is a linear sum assignment method that greedily performs bipartite graph matching to minimize costs, while Beam employs a best-first strategy, expanding partial solutions with the lowest-cost matches at each step to approximate GED. Hungarian (Riesen and Bunke, 2009) is not included as it has been incorporated in GOTSim. We generally use the official implementations and hyperparameters provided by the authors, with the VJ and Beam implementations sourced from (Yang and Zou, 2021). For models predicting the normalized GED value, we remove their last sigmoid layer to ensure a fair comparison.
5.1.3. Experimental Protocol
We partitioned the datasets into training and test graphs at a 4:1 ratio, with 20% of the training graphs used as the validation graphs. Training graph pairs are derived from the training graphs. All models are trained on a NVIDIA A6000 GPU with AMD EPYC 7282 CPU. The batch size is set to 128 graph pairs for training, 2048 for validation and test. Each epoch comprises 100 iterations due to the abundance of graph pairs, except for synthetic graphs. Each model is trained for ten epochs as a warm-up phase and tested on the validation set every epoch. Early stopping is applied with a patience of 50, meaning training stops if the validation loss fails to decrease for 50 consecutive validation steps to prevent overfitting, while ensuring that each model is sufficiently trained.
5.1.4. Implementation Details
We apply a 5-layer Graph Isomorphism Network (GIN) (Xu et al., 2019) with jump connections as GNN encoder to neighborhood information, whose hidden state and output are of 64 dimensionality. All MLPs in this work have adopted layer normalization and applied ReLU as the activation function. The outputs of these MLPs are of a dimensionality of 32. We used Adam as the optimizer, with a fixed learning rate of 0.001.
5.1.5. Evaluation Metrics
We evaluate models based on prediction error and ranking ability. Prediction error measures the point-wise difference between predicted and ground-truth GED values. Ranking ability is assessed by simulating retrieval scenarios in graph databases, where training and validation graphs serve as data graphs and test graphs as queries. During testing, each model calculates the GED between the query and data graphs, returning the most relevant data graphs with the lowest predicted GED for each query. For prediction error, we use Root Mean Square Error (RMSE) and Mean Absolute Error (MAE). Similar to MSE, RMSE is sensitive to outliers but presents errors in the original units for better interpretability, while MAE captures general errors. For ranking performance, we use Spearman’s Rank Correlation Coefficient ($\rho$), Kendall’s Rank Correlation Coefficient ($\tau$), and Precision at $k$ ($p@k$). $\rho$ sensitive to the magnitude of the differences in ranks, while $\tau$ evaluates relative order of pairs. $p@k$ calculates the intersection between the predicted top $k$ data graphs and the ground-truth top $k$ results. We sort all data graphs for each query based on both the ground truth GEDs and predictions in ascending order and compare the rankings. We report the average and standard deviation of results across five runs.

### v2

@@ 3 Background
Total $\text{\emph{Total Cost}: }c(S,T)_{f}=\sum_{s\in\mathcal{V}_{S},t\in\mathcal{V}_{T}}\text{c}(s,t)+\sum_{e\in f(\mathcal{A}_{S})\setminus\mathcal{A}_{T}}\text{c}(e,\epsilon)+\sum_{e^{\prime}\in\mathcal{A}_{T}\setminus f(\mathcal{A}_{S})}\text{c}(\epsilon,e^{\prime});$ (2)
[…]
In Eq. (1), $\text{c}(i_{s},i_{t})$ denotes the cost of editing element $i_{s}$ into $i_{t}$ via operation $op_{(i_{s},i_{t})}$, where $\text{cost}(\cdot)$ is a non-negative, customizable cost setting that specifies the operation cost, and $d(i_{s},i_{t})$ quantifies the attribute or structural dissimilarity between elements, typically 0 or 1 for individual nodes or edges. This formulation explicitly highlights the influence of operation costs in determining optimal mappings. In Eq. (2), $\mathcal{A}_{S}$ and $\mathcal{A}_{T}$ encode the intra-graph connections within the original graphs, and $\epsilon$ denotes a nonexistent edge. The terms $e\in f(\mathcal{A}_{S})\setminus\mathcal{A}_{T}$ and $e^{\prime}\in\mathcal{A}_{T}\setminus f(\mathcal{A}_{S})$ represent edges whose existence is not preserved under the mapping $f$, capturing the inter-graph discrepancies. Together, Eq. (2) models the graph-level dependencies between matches, reflecting how node correspondences jointly affect global structural alignment. Finally, in Eq. (3), $\text{GED}(S,T)$ is defined as the minimum of such total costs over all possible mappings.
4 Analyzing Flexible GED Computation in Embedding Space
Given graph pairs as input, learning-based node-level comparison methods approximate GED in the embedding space. They generally have three stages: (1) Encoding stage: they project the local structure of node onto the embedding space with an injective function $\Phi:\mathbf{X}\mapsto\mathbb{E}^{|\mathcal{V}|\times d}$, where $d$ is the dimensionality of the embeddings, $|\mathcal{V}|$ is the cardinality of $\mathcal{V}$; (2) Scoring stage: based on a distance/similarity measure function $\Psi$, they evaluate the distance for all possible source-target node embeddings; (3) Aligning stage: they use the pairwise embedding distances to guide the selection of matches $\mathbf{1}$, then use the total cost of a learned mapping to approximate GED.
In this work, we focus on the Scoring stage and the Aligning stage, as they are directly responsible for flexible and efficient GED computation. Existing methods typically apply cost-agnostic distance functions in the Scoring stage and use isolated node-level distances to guide the mapping establishment in the Aligning stage. However, these practices overlook the fundamental GED computation principles, namely operation cost sensitivity and graph-level dependencies between matches. Consequently, the resulting mappings are static and often suboptimal, necessitating costly post hoc refinements.
Limitation of Cost-agnostic Distance. Existing methods either ignore operation costs or consider them only at the end of the Aligning stage, treating them as irrelevant to the determination of optimal mappings. This results in static mappings that capture only intrinsic element dissimilarities. This limitation can be theoretically revealed by Eq. (1): When the cost setting function $\text{cost}(\cdot)$ assigns uniform operation cost (e.g., 1) to all operations, the total cost of a mapping corresponds to the number of edits. Let $f^{*}$ denote the optimal mapping under this uniform setting. When $\text{cost}(\cdot)$ varies, the total cost of $c(S,T)_{f^{*}}$ also changes, often rendering $f^{*}$ suboptimal, as shown in Figure 1. Consequently, existing methods cannot adapt to varying cost settings, limiting both the theoretical expressiveness and practical applicability of learned graph similarity models.
Limitation of Node-level Guidance. With the edit cost of pairwise node embeddings, the total cost of a learned mapping is given by:
$c(S,T)=\sum_{s\in\mathcal{V}_{S},t\in\mathcal{V}_{T}}\mathbf{1}_{s,t}\cdot c(s,t).$ (4)
Figure 2: Models marked with ‘*’ perform only a single round of refinement.
Where $\mathbf{1}_{s,t}$ denotes the (soft) selection of node correspondences. Existing methods typically determine $\mathbf{1}_{s,t}$ based on node-level distances. A representative example is GMN [16]. It derives local attention weights from pairwise node distances to guide inter-graph node representation fusions, and then refines node representations based on graph structures. Another representative approach is GraphEdX [13], which optimizes soft mappings via Gumbel-Sinkhorn networks, minimizing transport cost defined over node-level distances in a matrix form.
Both methods follow what we term reactive refinement strategies. They passively refine mappings based on local similarity signals without considering the graph-level dependencies between matches. This practice often leads to suboptimal intermediate mappings and costly refinements. The detailed comparison of methods can be found in Appendix A. Figure 2 provides empirical evidence that, when the number of refinement iterations in these models is limited to one, their MAE increases by 15% to 89% compared with the original settings recommended by the authors. This performance drop highlights their reliance on costly iterative refinements. We argue that efficient node-level GED approximation requires a proactive matching strategy that adheres to the principles of GED computation. Such a strategy should explicitly account for graph-level dependencies from the outset, capturing the joint impact of node matches on global structural alignment, thereby improving alignment quality and reducing computational overhead.
5 The Proposed GEN Architecture
Figure 3: Overview of Gen. The dashed node represents the dummy node.
An overview of the proposed Gen framework is illustrated in Figure 3. Given a source-target graph pair and customized costs per operation type for each possible match as inputs, Gen first uses a shared GNN Encoder as $\Phi$ to embed both graphs into a unified representation space, ensuring semantic alignment between node embeddings. Building on the analysis in Section 4, Gen introduces two components: the Cost-sensitive Scoring module $\Psi$ and the Proactive Guidance Optimization strategy, which jointly guide the soft selection of matches $\mathbf{1}$. $\Psi$ dynamically integrates operation costs with embedding distances, enabling adaptation to varying cost settings. The second module proactively optimizes guidance to capture the graph-level matching dependencies. It begins by estimating each node’s alignment difficulty to ensure global comparability. Then it propagates these difficulties within and across graphs via a Difficulty Propagation mechanism to capture intra-graph connections and inter-graph discrepancies. Gen uses the resulting guidance to construct $\mathbf{1}$, enabling informed one-shot matching and minimizing the need for reactive refinements. Finally, Gen computes the total cost by summing the weighted edit costs. The pseudo code of Gen is available in Appendix B.
5.1 Cost-sensitive Scoring
The node representation for node $v$ is initialized as the one-hot encoding of the node label, $x_{v}\in\mathbb{R}^{|\mathcal{X}|}$, which $|\mathcal{X}|$ is the number of label types. Gen first employs a $k$-layer GNN as $\Phi$ to generate node representations $\mathbf{X}^{k}$ based on the $k$-hop local structures. Due to the space limitation, the details of node feature initialization and the encoding process of $\Phi$ are provided in Appendix B.1. Given the initial label encoding $\mathbf{X}^{0}$ and node representations $\mathbf{X}^{k}$, this module computes the edit cost between all possible node pairs. For each node $v$, Gen first concatenate its $\mathbf{x}_{v}^{0}$ and $\mathbf{x}_{v}^{k}$: $\mathbf{h}_{v}=\text{Concat}(\mathbf{x}_{v}^{0},\mathbf{x}_{v}^{k})$, to emphasize label mismatches, where $\mathbf{h}_{v}\in\mathbb{R}^{|\mathcal{X}|+d}$.
Following Eq. (1), the edit cost from $s\in\mathcal{V}_{S}$ to $t\in\mathcal{V}_{T}$ is the operation cost-weighted inherent distance between $s$ and $t$ in the embedding space. We compute the inherent distance as $\text{d}(s,t)=\mathbf{h}_{s}-\mathbf{h}_{t}$. Instead of calculating distance involved in each operation type, Gen simplifies computation by using a unified $\text{d}(s,t)$ that represents all elements that should be edited. Given the predefined operation costs $cost{(s,t)}\in\mathbb{R}^{|op|}$ for each possible match, the edit cost is estimated as:
$c(s,t)=\text{Estimator}\big{(}\text{Concat}(cost{(s,t)},d(s,t))\big{)},$ (5)
where Estimator is a trainable multilayer perceptron (MLP). This design avoids naïvely applying each operation cost to the entire distance vector, which may contain elements unrelated to certain operations. By leveraging fully connected layers, the MLP can learn how different operation types interact with the latent differences in embeddings and produce a cost-aware correction.
Notably, Gen is the first method to incorporate operation costs before aligning, ensuring that while the inherent distances between embeddings remain consistent, they can be dynamically adjusted to accommodate varying cost settings. This is theoretically aligned with the GED computation principle in Eq. (1), enhancing the model’s understanding of GED computation. Further validation is provided in Experiment 2, highlighting the limitations of existing methods on this issue and the effectiveness of our solution.
5.2 Proactive Guidance Optimization
Unlike previous work that relies on the insufficient node-level distance as guidance, and reactively refines the match selections $\mathbf{1}$. Gen proactively optimize the guidance to capture the graph-level matching dependencies, constructing $\mathbf{1}$ in a single step.
@@ 5.2 Proactive Guidance Optimization
Guidance Initialization. Edit costs between embeddings can vary significantly across matches. While the absolute cost of each match may be clear, comparing these costs directly can be misleading, as some matches might seem optimal for one node but not contribute optimally to the overall graph alignment, leading to myopic, locally optimal decisions for individual nodes. To address this, Gen proposes constructing global comparability. It first normalizes the edit costs into global cost ratios:
$\gamma(s,t)=\frac{\exp(c(s,t))}{\sum_{s\in\mathcal{V}_{S},t\in\mathcal{V}_{T}}\exp(c(s,t))},$ (6)
@@
where $\gamma(s,t)\in\mathbb{R}^{d}$ represents the cost ratio of a match compared with the overall cost of all candidate matches. Intuitively, a match with a higher matching cost can have a larger ratio, reflecting its global significance within the graph-level alignment. Each node is involved in different candidate matches, forming a spectrum of differences between the node and the other graph. Gen gathers the cost ratios for each involved node to estimate the node alignment difficulty.
$\tilde{\gamma}(s)$ $=\sum_{i\in\mathcal{V}_{T}}\gamma(s,i),\quad\tilde{\gamma}(t)=\sum_{i\in\mathcal{V}_{S}}\gamma(i,t).$ (7)
@@
A higher alignment difficulty $\tilde{\gamma}(\cdot)$ indicates that accurate matching of the corresponding node is crucial to minimizing overall graph edit cost, providing a comprehensive assessment of each node’s relative importance.
Difficulty Propagation. Matching a node constrains the feasible matches of others, as neighborhood consistency must be preserved to minimize structural differences. Thus, the alignment difficulty of a node not only reflects its own ambiguity but also influences its neighbors. To capture such intra-graph connections, Gen propagates alignment difficulties within each graph:
$\hat{\gamma}(\mathcal{N}(v))=\text{MLP}\bigg{(}\text{Sum}\big{(}\tilde{\gamma}(v),\sum_{\mathcal{A}(v,u)=1}\tilde{\gamma}(u)\big{)}\bigg{)},$ (8)
where $\mathcal{N}(v)$ denotes the neighborhood of $v$. To strengthen the connection between difficulties and representations, Gen further refines $\hat{\gamma}(\mathcal{N}(v))$ using $\mathbf{h}_{v}$ as residual information:
$\gamma(\mathcal{N}(v))=\text{MLP}\bigg{(}\text{Concat}\big{(}\hat{\gamma}(\mathcal{N}(v)),\mathbf{h}_{v}\big{)}\bigg{)}.$ (9)
@@
Here, $\gamma(\mathcal{N}(v))\in\mathbb{R}^{d}$ is the neighborhood alignment difficulty for each node $v\in(\mathcal{V}_{S}\cup\mathcal{V}_{T})$, broadening the decision-making scope from individual nodes to their local contexts.
@@
A match connects two nodes within different graphs, each node can have a distinct alignment difficulty. Finally, for each node pair $(s,t)$, Gen considers such inter-graph discrepancies to balancing the alignment difficulties of the involved nodes. It computes certainty scores to approximate $\mathbf{1}_{s,t}$:
$\mathbf{1}_{s,t}=\text{MLP}\left(\text{Concat}(\gamma(s),\gamma(t))\right),$ (10)
where $\mathbf{1}_{s,t}\in\mathbb{R}^{d}$ reflects the soft confidence in matching $s$ and $t$, enabling more informed decisions in a single step by considering the graph-level dependencies between matches. Notably, all steps in this strategy are efficiently implemented using scatter operations over the sparse graph structure, which are well-optimized for GPU parallelism.
5.3 Prediction and Training Objective
Prediction. Given the $\mathbf{1}_{s,t}$ for each match, the edit cost per node is computed as:
$c(s)=\sum_{i\in\mathcal{V}_{T}}\mathbf{1}_{s,i}\cdot c(s,i),\quad c(t)=\sum_{i\in\mathcal{V}_{S}}\mathbf{1}_{i,t}\cdot c(i,t).$ (11)
Finally, the total cost of mappings is calculated as the sum of certainty-weighted edit costs, according to Eq. (4). The total costs for mappings established for graphs $S$ and $T$ are:
$c(S,T)_{f_{S}}=\sum_{s\in\mathcal{V}_{S}}c(s),\quad c(S,T)_{f_{T}}=\sum_{t\in\mathcal{V}_{T}}c(t).$ (12)
The predicted GED is obtained as:
$\text{GED}(S,T)=\text{Linear}\left(\text{Mean}\left(c(S,T)_{f_{S}},c(S,T)_{f_{T}}\right)\right),$ (13)
@@ 5.3 Prediction and Training Objective
where $\text{Linear}(\cdot)$ reduces the dimensionality from $d$ to $1$. This process indirectly encourages consistent mappings between the $S$ and $T$, or the average matching cost increases accordingly.
Training Objective. Given the predicted GED $\hat{y}$ and the ground-truth GED $y$, Gen minimizes the Huber loss [12], which is computed as follows.
@@
$\mathcal{L}(\hat{y},y)=\begin{cases}\frac{1}{2}(y-\hat{y})^{2}&\text{if }|y-\hat{y}|\leq\delta,\\ \delta\left(|y-\hat{y}|-\frac{1}{2}\delta\right)&\text{otherwise},\end{cases}$ (14)
where $\delta$ controls the transition between quadratic and linear behaviors, ensuring stable training under diverse error scales.
5.4 Complexity Analysis
The complexity of the GNN Encoder is $\mathcal{O}(k|\mathcal{E}|d_{\text{hidden}})$, where $k$ is the number of GNN layers and $\mathcal{E}$ is the set of edges. The complexity of Cost-sensitive Scoring is $\mathcal{O}(\max(|\mathcal{V}_{S}|,|\mathcal{V}_{T}|)^{2}\cdot d)$. The Proactive Optimization Strategy introduces an additional complexity of $\mathcal{O}(2(\max(|\mathcal{V}_{S}|,|\mathcal{V}_{T}|)^{2}d+|\mathcal{E}|d))$.